A 110-V electric iron draws 3.0 A of current. How much thermal energy is developed each hour?
Answer #1

V = IR

so R = V / I =110V / 3.0A = 37 \Omega

and Q = E = I2Rt = (3.0 A)2(37 \Omega)(1 h)(3600s/h) = 1.2 x 106 J

luck !
Answer #2

ohms law, here v is the voltage, l is the current, R is the resistance び110び 1 3.0A R 36.660 Thermal energy. E-12 Rt-G.0AY *